OpenStudy (anonymous):

if y`=(9x^4+48x^3+37x^2-20x+3)/(9x^3+12x^2-11x+2) find y y`means the derivative of y please help

OpenStudy (anonymous):

I think you need to find the antiderivative.

OpenStudy (anonymous):

i tried i used division but i couldn`t simplify the denominator

OpenStudy (anonymous):

Subtract the power by 1 and the denominator by one also.

OpenStudy (anonymous):

x^4. = x^4-1 4-1

OpenStudy (anonymous):

So you get 1/3x^3

OpenStudy (anonymous):

9(1/3)= 3 So you get for the first one 3x^3 Hmmmmmm that doesn't seen to be working.

OpenStudy (anonymous):

Do you have the answer? Is the first one 9/5x^5?

OpenStudy (anonymous):

If it is instead of subtracting 1 you need to add 1 to what I did above.

OpenStudy (anonymous):

impartial fractions maybe

OpenStudy (anonymous):

partial fractions my bad

OpenStudy (anonymous):

X^(n+1)/n+1

OpenStudy (anonymous):

i don`t have the answer yeah >>>partial fractions i`ii try the(n+1) method thanx

OpenStudy (anonymous):

use long division to break down the fraction cause num.power has to be smaller than denom,

OpenStudy (anonymous):

i did the division couldn`t simplify the denominator

OpenStudy (anonymous):

9/5x^5+48/4x^4+37/3x^3-20/2x^2+3x